A sphere of radius R carries a charge of volume charge density ρ= ar, where a is a


constant and r denotes the distance from the centre of the sphere. Calculate the


total charge enclosed by the sphere and the electric field at points lying inside and


outside the sphere.

1
Expert's answer
2022-02-28T10:41:46-0500

"\rho=ar,"

"E(4\pi r^2)=\frac{\int\rho\cdot4\pi r^2dr}{\varepsilon_0}=\frac{\int ar\cdot4\pi r^2dr}{\varepsilon_0},"

"E=\frac{ar^2}{4\varepsilon_0}."